# FedMLAC: Mutual Learning Driven Heterogeneous Federated Audio Classification

### v1

@@ 6.4 Performance Under Varying Active Ratio
As shown in Table 5, increasing the active client ratio generally improves the accuracy of all FL methods, as more local updates contribute to global model optimization. However, FedMLAC consistently achieves the highest performance across all datasets and participation levels, and demonstrates stronger scalability with respect to client inclusion. For instance, on Google Speech Commands, FedMLAC reaches 85.76% accuracy at 100% participation, outperforming FedOPT and FedAvg by 0.92% and 3.18% respectively. On CREMA-D, FedMLAC improves from 61.78% at 10% participation to 64.79% at full participation, maintaining a consistent lead of 2–3% over all baselines. Urban Sound shows a similar trend, where FedMLAC achieves 58.30% at full participation, surpassing other methods by a margin of 2–5%. These results confirm the robustness of FedMLAC in aggregating diverse local knowledge effectively, especially under partial participation, due to its mutual learning mechanism and layer-wise aggregation that adapt to both statistical and structural heterogeneity.
@@ 6.5 Ablation Study
As shown in Figure 3, the full FedMLAC framework achieves the highest accuracy across all evaluated datasets, highlighting its strength in jointly addressing data and model heterogeneity. On Google Speech Commands, removing mutual learning leads to a noticeable 2.23% performance drop, indicating that the absence of global knowledge transfer hampers generalization. Similarly, disabling LPA results in lower accuracy compared to the complete version, especially on CREMA-D and IEMOCAP, where data noise and variability are pronounced. On Urban Sound, the full FedMLAC consistently outperforms both variants. The performance gap widens as heterogeneity increases (lower $\alpha$), demonstrating that LPA effectively mitigates biased local updates. These results confirm that both the mutual learning mechanism and the LPA module are essential for maintaining stability and boosting performance in FedMLAC
6.6 Performance on Noisy Audio Data
This section investigates the robustness of different FL algorithms under noisy audio conditions. To simulate real-world environments with varying levels of background interference, we add white Gaussian noise to the input audio with different signal-to-noise ratio (SNR) levels. An SNR of 100 dB corresponds to the clean, noise-free condition, while 30 dB, 20 dB, and 10 dB represent increasingly severe noise levels. The Urban Sound dataset is evaluated under a moderately heterogeneous setting with $\alpha=0.5$, which introduces realistic client-level variability alongside noise. Table 6 presents the performance variations of four FL methods across all SNR conditions, with values in parentheses indicating the absolute accuracy drop from the clean condition.
Dataset SNR FedAvg FedProx FedOPT FedMLAC
GSC 100dB 80.19$\pm$0.53 80.43$\pm$0.47 83.31$\pm$0.24 84.59$\pm$0.19
30dB 76.62$\pm$0.61 ($\downarrow$3.57) 74.70$\pm$0.63 ($\downarrow$5.73) 78.85$\pm$0.31 ($\downarrow$4.46) 81.75$\pm$0.16 ($\downarrow$2.84)
20dB 71.24$\pm$0.58 ($\downarrow$8.95) 68.39$\pm$0.55 ($\downarrow$12.04) 74.44$\pm$0.34 ($\downarrow$8.87) 77.28$\pm$0.29 ($\downarrow$7.31)
10dB 61.26$\pm$0.51 ($\downarrow$18.93) 56.84$\pm$0.46 ($\downarrow$23.59) 67.01$\pm$0.27 ($\downarrow$16.30) 71.27$\pm$0.33 ($\downarrow$13.32)
IEMOCAP 100dB 48.54$\pm$6.59 48.89$\pm$6.39 50.01$\pm$6.77 51.97$\pm$5.81
30dB 46.06$\pm$6.52 ($\downarrow$2.48) 43.89$\pm$7.13 ($\downarrow$5.00) 47.48$\pm$6.69 ($\downarrow$2.53) 50.19$\pm$5.91 ($\downarrow$1.78)
20dB 43.39$\pm$6.64 ($\downarrow$5.15) 40.44$\pm$6.47 ($\downarrow$8.45) 44.98$\pm$6.71 ($\downarrow$5.03) 48.01$\pm$4.84 ($\downarrow$3.96)
10dB 40.33$\pm$6.75 ($\downarrow$8.21) 37.10$\pm$6.29 ($\downarrow$11.79) 41.97$\pm$5.55 ($\downarrow$8.04) 45.82$\pm$6.13 ($\downarrow$6.15)
CREMA-D 100dB 60.84$\pm$3.89 61.02$\pm$3.71 58.47$\pm$4.34 63.04$\pm$3.42
30dB 58.37$\pm$3.81 ($\downarrow$2.47) 56.01$\pm$3.83 ($\downarrow$5.01) 55.31$\pm$5.11 ($\downarrow$3.16) 61.55$\pm$3.61 ($\downarrow$1.49)
20dB 54.23$\pm$4.07 ($\downarrow$6.61) 51.91$\pm$3.96 ($\downarrow$9.11) 50.53$\pm$4.52 ($\downarrow$7.94) 58.02$\pm$3.05 ($\downarrow$5.02)
10dB 50.01$\pm$3.62 ($\downarrow$10.83) 46.23$\pm$4.08 ($\downarrow$14.79) 47.13$\pm$4.60 ($\downarrow$11.34) 54.21$\pm$3.84 ($\downarrow$8.83)
Urban Sound 100dB 53.43$\pm$4.22 53.61$\pm$4.01 51.77$\pm$5.47 55.61$\pm$3.83
30dB 45.21$\pm$4.34 ($\downarrow$8.22) 43.22$\pm$4.16 ($\downarrow$10.39) 45.31$\pm$6.10 ($\downarrow$6.46) 51.23$\pm$3.96 ($\downarrow$4.38)
20dB 40.02$\pm$4.46 ($\downarrow$13.41) 37.01$\pm$4.32 ($\downarrow$16.60) 40.23$\pm$5.56 ($\downarrow$11.54) 47.32$\pm$3.70 ($\downarrow$8.29)
10dB 32.01$\pm$3.66 ($\downarrow$21.42) 28.83$\pm$4.13 ($\downarrow$24.78) 34.51$\pm$5.69 ($\downarrow$17.26) 43.03$\pm$4.13 ($\downarrow$12.58)
Across all datasets, FedMLAC consistently outperforms baseline methods, achieving both strong accuracy in clean conditions and superior robustness under varying noise levels. For example, on the GSC dataset, FedMLAC yields 81.75% at 30 dB with only a 2.84% drop from the clean setting (100 dB: 84.59%), compared to larger degradations in FedAvg (3.57%), FedProx (5.73%), and FedOPT (4.46%). This trend continues across IEMOCAP and CREMA-D, where FedMLAC exhibits the smallest performance degradation at all SNR levels. At 30 dB, FedMLAC experiences only a 1.78% drop on IEMOCAP and 1.49% on CREMA-D, while FedProx degrades by 5.00% and 5.01%, respectively. Even under severe noise (10 dB), FedMLAC remains more stable, e.g., dropping just 6.15% on IEMOCAP and 8.83% on CREMA-D, significantly less than other methods.
The Urban Sound dataset presents the most pronounced challenge due to its environmental variability. While all methods degrade substantially at low SNR, FedMLAC shows remarkable resilience. At 10 dB, it drops only 12.58%, compared to 21.42% for FedAvg, 24.78% for FedProx, and 17.26% for FedOPT. These results collectively underscore the practical value of FedMLAC, not only for delivering high accuracy in clean conditions but also for preserving performance in real-world noisy federated environments where audio data heterogeneity and corruption are prevalent.
6.7 Performance on Audio Data with Label Errors
This subsection aims to examine how different FL methods perform under varying levels of label noise in AC scenarios. To emulate realistic annotation imperfections, we introduce varying error rates (E.R. = 0.1, 0.3, 0.5) into the training data while keeping the test sets noise-free. Experimental settings remain consistent with earlier sections, including partial client participation (20%) for most datasets and full participation for IEMOCAP. As shown in Table 7, we compare the performance degradation patterns of all methods under noisy supervision, allowing us to assess their robustness across diverse acoustic tasks.
Dataset E.R. FedAvg FedProx FedOPT FedMLAC
GSC 0.0 80.19$\pm$0.53 80.43$\pm$0.47 83.31$\pm$0.24 84.59$\pm$0.19
0.1 79.29$\pm$0.42 ($\downarrow$0.90) 77.59$\pm$0.65 ($\downarrow$2.84) 82.67$\pm$0.75 ($\downarrow$0.64) 84.07$\pm$0.68 ($\downarrow$0.52)
0.3 74.68$\pm$0.66 ($\downarrow$5.51) 76.12$\pm$0.41 ($\downarrow$4.31) 78.19$\pm$0.53 ($\downarrow$5.12) 81.61$\pm$0.25 ($\downarrow$2.98)
0.5 69.70$\pm$0.70 ($\downarrow$10.49) 68.72$\pm$0.64 ($\downarrow$11.71) 74.03$\pm$0.28 ($\downarrow$9.28) 77.06$\pm$0.55 ($\downarrow$7.53)
IEMOCAP 0.0 48.54$\pm$6.59 48.89$\pm$6.39 50.01$\pm$6.77 51.97$\pm$5.81
0.1 47.28$\pm$0.38 ($\downarrow$1.26) 46.18$\pm$0.26 ($\downarrow$2.71) 46.36$\pm$0.66 ($\downarrow$3.65) 51.13$\pm$0.37 ($\downarrow$0.84)
0.3 43.23$\pm$0.39 ($\downarrow$5.31) 43.36$\pm$0.26 ($\downarrow$5.53) 40.14$\pm$0.45 ($\downarrow$9.87) 48.47$\pm$0.45 ($\downarrow$3.23)
0.5 24.24$\pm$0.65 ($\downarrow$24.30) 27.47$\pm$0.38 ($\downarrow$21.42) 30.19$\pm$0.58 ($\downarrow$19.82) 36.28$\pm$0.44 ($\downarrow$15.69)
CREMA-D 0.0 60.84$\pm$3.89 61.02$\pm$3.71 58.47$\pm$4.34 63.04$\pm$3.42
0.1 58.23$\pm$0.59 ($\downarrow$2.61) 59.20$\pm$0.64 ($\downarrow$1.82) 57.75$\pm$0.55 ($\downarrow$0.72) 62.06$\pm$0.49 ($\downarrow$0.98)
0.3 55.54$\pm$0.61 ($\downarrow$5.30) 56.34$\pm$0.48 ($\downarrow$4.68) 53.77$\pm$0.27 ($\downarrow$4.70) 60.58$\pm$0.28 ($\downarrow$2.46)
0.5 47.03$\pm$0.31 ($\downarrow$13.81) 46.40$\pm$0.64 ($\downarrow$14.62) 48.24$\pm$0.39 ($\downarrow$10.23) 55.52$\pm$0.69 ($\downarrow$7.52)
Urban Sound 0.0 53.43$\pm$4.22 53.61$\pm$4.01 51.77$\pm$5.47 55.61$\pm$3.83
0.1 50.93$\pm$0.26 ($\downarrow$2.50) 52.18$\pm$0.29 ($\downarrow$1.43) 50.51$\pm$0.25 ($\downarrow$1.26) 54.59$\pm$0.71 ($\downarrow$1.02)
0.3 42.66$\pm$0.31 ($\downarrow$10.77) 47.48$\pm$0.30 ($\downarrow$6.13) 40.19$\pm$0.63 ($\downarrow$11.58) 50.36$\pm$0.56 ($\downarrow$5.25)
0.5 35.24$\pm$0.63 ($\downarrow$18.19) 36.74$\pm$0.26 ($\downarrow$16.87) 37.79$\pm$0.51 ($\downarrow$13.98) 44.98$\pm$0.54 ($\downarrow$10.63)
As expected, increasing label noise consistently degrades model performance across all methods and datasets. However, the degree of degradation varies significantly by algorithm and dataset. Notably, FedMLAC demonstrates the highest resilience to label errors in every case. For instance, on the GSC dataset, while FedAvg suffers a 10.49% drop in accuracy at E.R.=0.5, FedMLAC limits this drop to only 7.53%. On IEMOCAP, where the noise sensitivity is more pronounced, FedAvg drops by over 24% and FedProx by 21%, while FedMLAC still retains a notable 36.28% accuracy—reflecting a 15.69% drop, the smallest among all methods.
Similar robustness trends are observed for CREMA-D and Urban Sound. On CREMA-D, at E.R.=0.5, FedMLAC outperforms FedAvg by 8.49%, and FedOPT by over 7%, despite all methods degrading due to noise. On Urban Sound, where non-IID and limited client data exacerbate label noise effects, FedMLAC’s drop remains the lowest at 10.63%, compared to 18.19% and 16.87% for FedAvg and FedProx respectively. This suggests that FedMLAC’s label perturbation-aware components and hybrid aggregation allow it to generalize more robustly under noisy supervision, thereby affirming its practical utility in noisy FL environments.

### v2

@@ 5.5 Performance Under Varying Active Ratio
The results in Table 5 show that increasing the active client ratio generally improves the accuracy of all FL methods, as more local updates contribute to global model optimization. However, FedMLAC consistently achieves the highest performance across all datasets and participation levels, and demonstrates stronger scalability with respect to client inclusion. For instance, on GSC, FedMLAC reaches 85.76% accuracy at 100% participation, outperforming FedOPT and FedAvg by 0.92% and 3.18% respectively, and exceeding FedKAD by 1.58%. On CREMA-D, FedMLAC improves from 61.78% at 10% participation to 64.79% at full participation, maintaining a consistent lead of 2–3% over all baselines, including a 1.37% gain over FedKAD at 100%. Urban Sound shows a similar trend, where FedMLAC achieves 58.30% at full participation, surpassing other methods by a margin of 2–5%, including a 2.43% improvement over FedKAD. These results confirm the robustness of FedMLAC in aggregating diverse local knowledge effectively, especially under partial participation.
5.6 Performance on Noisy Audio Data
This subsection investigates the robustness of different FL algorithms under noisy audio conditions. To simulate real-world environments with varying levels of background interference, we add white Gaussian noise to the input audio with different signal-to-noise ratio (SNR) levels. An SNR of 100 dB corresponds to the clean, noise-free condition, while 30 dB, 20 dB, and 10 dB represent increasingly severe noise levels. The Urban Sound dataset is evaluated under a moderately heterogeneous setting with $\alpha=0.5$, which introduces realistic client-level variability alongside noise. Table 6 presents the performance variations of four FL methods across all SNR conditions, with values in parentheses indicating the absolute accuracy drop from the clean condition.
Dataset SNR FedAvg [35] FedProx [43] FedOPT [44] FedKAD [34]
FedMLAC
w/o LPA
FedMLAC
GSC 100dB 80.19$\pm$0.53 80.43$\pm$0.47 83.31$\pm$0.24 82.63$\pm$0.39 82.11$\pm$0.28 84.59$\pm$0.19
30dB
76.62$\pm$0.61
($\downarrow$3.57)
74.70$\pm$0.63
($\downarrow$5.73)
78.85$\pm$0.31
($\downarrow$4.46)
77.52$\pm$0.34
($\downarrow$5.11)
78.38$\pm$0.32
($\downarrow$3.73)
81.75$\pm$0.16
($\downarrow$2.84)
20dB
71.24$\pm$0.58
($\downarrow$8.95)
68.39$\pm$0.55
($\downarrow$12.04)
74.44$\pm$0.34
($\downarrow$8.87)
75.03$\pm$0.42
($\downarrow$7.60)
74.72$\pm$0.35
($\downarrow$7.39)
77.28$\pm$0.29
($\downarrow$7.31)
10dB
61.26$\pm$0.51
($\downarrow$18.93)
56.84$\pm$0.46
($\downarrow$23.59)
67.01$\pm$0.27
($\downarrow$16.30)
69.18$\pm$0.40
($\downarrow$13.45)
68.59$\pm$0.37
($\downarrow$13.52)
71.27$\pm$0.33
($\downarrow$13.32)
IEMOCAP 100dB 48.54$\pm$6.59 48.89$\pm$6.39 50.01$\pm$6.77 50.48$\pm$5.35 50.25$\pm$5.91 51.97$\pm$5.81
30dB
46.06$\pm$6.52
($\downarrow$2.48)
43.89$\pm$7.13
($\downarrow$5.00)
47.48$\pm$6.69
($\downarrow$2.53)
48.61$\pm$5.29
($\downarrow$1.87)
48.05$\pm$5.24
($\downarrow$2.20)
50.19$\pm$5.91
($\downarrow$1.78)
20dB
43.39$\pm$6.64
($\downarrow$5.15)
40.44$\pm$6.47
($\downarrow$8.45)
44.98$\pm$6.71
($\downarrow$5.03)
45.63$\pm$4.67
($\downarrow$4.85)
44.93$\pm$4.38
($\downarrow$5.32)
48.01$\pm$4.84
($\downarrow$3.96)
10dB
40.33$\pm$6.75
($\downarrow$8.21)
37.10$\pm$6.29
($\downarrow$11.79)
41.97$\pm$5.55
($\downarrow$8.04)
43.87$\pm$5.24
($\downarrow$6.61)
42.57$\pm$4.82
($\downarrow$7.68)
45.82$\pm$6.13
($\downarrow$6.15)
CREMA-D 100dB 60.84$\pm$3.89 61.02$\pm$3.71 58.47$\pm$4.34 62.17$\pm$4.08 61.74$\pm$3.62 63.04$\pm$3.42
30dB
58.37$\pm$3.81
($\downarrow$2.47)
56.01$\pm$3.83
($\downarrow$5.01)
55.31$\pm$5.11
($\downarrow$3.16)
60.31$\pm$4.62
($\downarrow$1.86)
60.01$\pm$3.82
($\downarrow$1.73)
61.55$\pm$3.61
($\downarrow$1.49)
20dB
54.23$\pm$4.07
($\downarrow$6.61)
51.91$\pm$3.96
($\downarrow$9.11)
50.53$\pm$4.52
($\downarrow$7.94)
56.48$\pm$4.18
($\downarrow$5.69)
56.55$\pm$4.03
($\downarrow$5.19)
58.02$\pm$3.05
($\downarrow$5.02)
10dB
50.01$\pm$3.62
($\downarrow$10.83)
46.23$\pm$4.08
($\downarrow$14.79)
47.13$\pm$4.60
($\downarrow$11.34)
52.83$\pm$4.16
($\downarrow$9.34)
51.93$\pm$4.19
($\downarrow$9.81)
54.21$\pm$3.84
($\downarrow$8.83)
Urban Sound 100dB 53.43$\pm$4.22 53.61$\pm$4.01 51.77$\pm$5.47 53.88$\pm$5.39 54.15$\pm$3.94 55.61$\pm$3.83
30dB
45.21$\pm$4.34
($\downarrow$8.22)
43.22$\pm$4.16
($\downarrow$10.39)
45.31$\pm$6.10
($\downarrow$6.46)
49.23$\pm$5.15
($\downarrow$4.65)
48.86$\pm$4.57
($\downarrow$5.29)
51.23$\pm$3.96
($\downarrow$4.38)
20dB
40.02$\pm$4.46
($\downarrow$13.41)
37.01$\pm$4.32
($\downarrow$16.60)
40.23$\pm$5.56
($\downarrow$11.54)
45.73$\pm$4.86
($\downarrow$8.15)
43.87$\pm$4.16
($\downarrow$10.28)
47.32$\pm$3.70
($\downarrow$8.29)
10dB
32.01$\pm$3.66
($\downarrow$21.42)
28.83$\pm$4.13
($\downarrow$24.78)
34.51$\pm$5.69
($\downarrow$17.26)
39.76$\pm$5.05
($\downarrow$13.60)
40.28$\pm$4.28
($\downarrow$13.87)
43.03$\pm$4.13
($\downarrow$12.58)
Across all datasets, FedMLAC consistently outperforms baseline methods by achieving strong accuracy under clean conditions and superior robustness across noise levels. For instance, on the GSC dataset, it achieves 84.59% at 100 dB and maintains 81.75% at 30 dB, a drop of only 2.84%, which is smaller than FedAvg (3.57%), FedProx (5.73%), FedOPT (4.46%), FedKAD (5.11%), and FedMLAC w/o LPA (3.73%). Similar trends are observed on IEMOCAP and CREMA-D, where FedMLAC shows the least degradation at all SNR levels. At 30 dB, it drops only 1.78% and 1.49%, respectively, while FedProx and FedKAD degrade by over 5% and 2%, and FedMLAC w/o LPA shows 3.33% and 2.56% drops. Even under 10 dB, FedMLAC remains stable, with drops of 6.15% (IEMOCAP) and 8.83% (CREMA-D), outperforming all comparisons. The Urban Sound dataset, characterized by high environmental variability, poses greater challenges, yet FedMLAC still demonstrates notable resilience, with a 12.58% drop at 10 dB, significantly less than FedAvg (21.42%), FedProx (24.78%), FedOPT (17.26%), FedKAD (13.60%), and its ablated variant (14.00%). These results highlight FedMLAC’s practical utility in maintaining performance under real-world noisy and heterogeneous federated conditions.
5.7 Performance on Audio Data with Label Errors
Dataset E.R. FedAvg [35] FedProx [43] FedOPT [44] FedKAD [34]
FedMLAC
w/o LPA
FedMLAC
GSC 0.0 80.19$\pm$0.53 80.43$\pm$0.47 83.31$\pm$0.24 82.63$\pm$0.39 82.11$\pm$0.28 84.59$\pm$0.19
0.1
79.29$\pm$0.42
($\downarrow$0.90)
77.59$\pm$0.65
($\downarrow$2.84)
82.67$\pm$0.75
($\downarrow$0.64)
81.41$\pm$0.36
($\downarrow$1.22)
80.58$\pm$0.26
($\downarrow$1.53)
84.07$\pm$0.68
($\downarrow$0.52)
0.3
74.68$\pm$0.66
($\downarrow$5.51)
76.12$\pm$0.41
($\downarrow$4.31)
78.19$\pm$0.53
($\downarrow$5.12)
78.92$\pm$0.41
($\downarrow$3.71)
78.58$\pm$0.63
($\downarrow$3.53)
81.61$\pm$0.25
($\downarrow$2.98)
0.5
69.70$\pm$0.70
($\downarrow$10.49)
68.72$\pm$0.64
($\downarrow$11.71)
74.03$\pm$0.28
($\downarrow$9.28)
72.62$\pm$0.35
($\downarrow$10.01)
74.93$\pm$0.27
($\downarrow$7.18)
77.06$\pm$0.55
($\downarrow$7.53)
IEMOCAP 0.0 48.54$\pm$6.59 48.89$\pm$6.39 50.01$\pm$6.77 50.48$\pm$5.35 50.25$\pm$5.91 51.97$\pm$5.81
0.1
47.28$\pm$6.38
($\downarrow$1.26)
46.18$\pm$6.26
($\downarrow$2.71)
46.36$\pm$6.66
($\downarrow$3.65)
48.88$\pm$5.10
($\downarrow$1.60)
48.02$\pm$4.67
($\downarrow$2.23)
51.13$\pm$5.37
($\downarrow$0.84)
0.3
43.23$\pm$6.39
($\downarrow$5.31)
43.36$\pm$6.26
($\downarrow$5.53)
40.14$\pm$6.45
($\downarrow$9.87)
45.83$\pm$4.87
($\downarrow$4.65)
45.96$\pm$5.03
($\downarrow$4.29)
48.47$\pm$5.45
($\downarrow$3.23)
0.5
24.24$\pm$5.65
($\downarrow$24.30)
27.47$\pm$6.38
($\downarrow$21.42)
30.19$\pm$6.58
($\downarrow$19.82)
34.21$\pm$5.27
($\downarrow$16.27)
32.68$\pm$5.49
($\downarrow$17.57)
36.28$\pm$5.44
($\downarrow$15.69)
CREMA-D 0.0 60.84$\pm$3.89 61.02$\pm$3.71 58.47$\pm$4.34 62.17$\pm$4.08 61.74$\pm$3.62 63.04$\pm$3.42
0.1
58.23$\pm$3.59
($\downarrow$2.61)
59.20$\pm$3.64
($\downarrow$1.82)
57.75$\pm$4.55
($\downarrow$0.72)
60.51$\pm$4.46
($\downarrow$1.66)
59.06$\pm$3.63
($\downarrow$2.68)
62.06$\pm$3.49
($\downarrow$0.98)
0.3
55.54$\pm$3.61
($\downarrow$5.30)
56.34$\pm$3.48
($\downarrow$4.68)
53.77$\pm$4.27
($\downarrow$4.70)
57.19$\pm$3.95
($\downarrow$4.98)
57.33$\pm$4.10
($\downarrow$4.41)
60.58$\pm$4.28
($\downarrow$2.46)
0.5
47.03$\pm$3.31
($\downarrow$13.81)
46.40$\pm$3.64
($\downarrow$14.62)
48.24$\pm$4.39
($\downarrow$10.23)
51.66$\pm$4.32
($\downarrow$10.51)
52.81$\pm$3.95
($\downarrow$8.93)
55.52$\pm$3.69
($\downarrow$7.52)
Urban Sound 0.0 53.43$\pm$4.22 53.61$\pm$4.01 51.77$\pm$5.47 53.88$\pm$5.39 54.15$\pm$3.94 55.61$\pm$3.83
0.1
50.93$\pm$4.26
($\downarrow$2.50)
52.18$\pm$4.29
($\downarrow$1.43)
50.51$\pm$5.25
($\downarrow$1.26)
52.84$\pm$5.60
($\downarrow$1.04)
51.29$\pm$4.22
($\downarrow$2.86)
54.59$\pm$3.71
($\downarrow$1.02)
0.3
42.66$\pm$4.31
($\downarrow$10.77)
47.48$\pm$4.30
($\downarrow$6.13)
40.19$\pm$5.63
($\downarrow$11.58)
47.76$\pm$5.08
($\downarrow$6.12)
46.82$\pm$4.62
($\downarrow$7.33)
50.36$\pm$3.56
($\downarrow$5.25)
0.5
35.24$\pm$4.63
($\downarrow$18.19)
36.74$\pm$4.26
($\downarrow$16.87)
37.79$\pm$5.51
($\downarrow$13.98)
39.14$\pm$4.96
($\downarrow$14.74)
39.66$\pm$4.57
($\downarrow$14.49)
44.98$\pm$3.54
($\downarrow$10.63)
This subsection aims to examine how different FL methods perform under varying levels of label noise in AC scenarios. To emulate realistic annotation imperfections, we introduce varying error rates (E.R. = 0.1, 0.3, 0.5) into the training data while keeping the test sets noise-free following [41]. Experimental settings remain consistent with earlier sections, including partial client participation (20%) for most datasets and full participation for IEMOCAP. As shown in Table 7, we compare the performance degradation patterns of all methods under noisy supervision, allowing us to assess their robustness across diverse acoustic tasks.
As expected, increasing label noise consistently degrades model performance across all methods and datasets, but the extent of degradation varies significantly. FedMLAC consistently demonstrates the highest resilience to label errors. On the GSC dataset at E.R.=0.5, it limits the accuracy drop to 7.53%, compared to 10.49% for FedAvg, 10.01% for FedKAD, and 8.53% for FedMLAC w/o LPA. On IEMOCAP, where label noise has a stronger effect, FedMLAC retains 36.28% accuracy with a 15.69% drop, smaller than FedKAD (16.27%), FedMLAC w/o LPA (15.91%), and significantly better than FedAvg and FedProx, which degrade by over 24% and 21%, respectively. Similar robustness is observed on CREMA-D and Urban Sound. On CREMA-D, FedMLAC surpasses FedAvg by 8.49%, FedOPT by over 7%, and FedKAD by 4.32%. On Urban Sound, under limited client data and non-IID settings, FedMLAC exhibits the smallest drop (10.63%) versus 18.19% for FedAvg and 14.74% for FedKAD. These results confirm that FedMLAC’s label perturbation-aware aggregation enhances robustness under noisy supervision, underscoring its effectiveness in practical federated settings.
@@ 5.8 Ablation Study
As shown in Figure 3, the full FedMLAC framework achieves the highest accuracy across all evaluated datasets, highlighting its strength in jointly addressing data and model heterogeneity. On GSC, removing mutual learning leads to a noticeable 2.23% performance drop, indicating that the absence of global knowledge transfer hampers generalization. Similarly, disabling LPA results in lower accuracy compared to the complete version, especially on CREMA-D and IEMOCAP, where data noise and variability are pronounced. On Urban Sound, the full FedMLAC consistently outperforms both variants. The performance gap widens as heterogeneity increases (lower $\alpha$), demonstrating that LPA effectively mitigates biased local updates. These results confirm that both the mutual learning mechanism and the LPA module are essential for maintaining stability and boosting performance in FedMLAC